# MotionFix: Text-Driven 3D Human Motion Editing

### v1

@@ 5.3 Ablations
Figure 4: TMED generations: We illustrate several generations from our model with overlaid source (red) and generated (blue) motions. We showcase a variety of test cases ranging from elaborate edits (top left) to high-level short commands such as “don’t get up” (row 4, column 1). TMED is able to perform both edits that describe temporal (e.g., “slow down”) or spatial (e.g., “raise your arms higher so it is overhead”) modifications.
Figure 5: Guidances of conditions: We illustrate the R@1 performance of TMED for generated-to-target (left) and generated-to-source (right) retrieval benchmarks for $s_{L},s_{M_{S}}\in[0,5]$.
Guidance hyperparameters. In Figure 5, we present how TMED performs across different guidance values for both conditions. $x$-axis controls the text guidance $s_{L}$, and $y$-axis controls the source motion guidance $s_{M_{S}}$ at test time. We report both generated-to-target (left) and source-to-target (right) retrieval results. We observe that increasing the source motion guidance leads to consistent improvement in R@1 for the generated-to-source retrieval across all $s_{L}$ values. This verifies that one can control the resemblance to the source motion by using $s_{M_{S}}$.
As per generated-to-target retrieval, although it appears to highly benefit from conditioning on the source motion, this does not always result in consistent performance improvements. Values of $s_{L}$ smaller than $2.0$ result in performance drop if $s_{M_{S}}$ is greater than $2.0$, which highlights the importance of the edit texts. However, for values greater than $2.0$, the increased conditioned on the source motion helps the generation to be more aligned with the target motions.
Figure 6: Qualitative comparisons with baselines: We provide example results by comparing TMED against the baselines on the MF test set: MDMS (top), MDM-BPS (middle) and MDM-BP (bottom). First two columns show the source (red) and ground-truth target (green) motions. Third column is reserved for baselines, the last column for TMED. Generations are denoted in blue.
5.4 Qualitative Results
We display several generations from TMED in Figure 4 to enable qualitative asssessment. We observe that our model can perform different types of edits such as the addition of actions (“rotate wrists instead of stretching like yawn”), temporal edits in a motion (“start by bending to the side”), speed edits (“sit little bit slower”, “slow down”) and combinations of these.
In Figure 6, we further show qualitative comparisons between TMED and various baselines. We provide two comparisons for each baseline (top block for MDMS, middle block for MDM-BPS, and bottom block for MDM-B).
@@ 5.4 Qualitative Results
Finally, we illustrate generations from our strongest baseline MDM-BP in the last two rows. Both generations involve all parts of the body (e.g., “slow down”) making it hard to follow the source motion. In comparison, our model faithfully performs most of the edits. The disadvantage of TMED, on the other hand, might be the generalization to motion pairs where the TMR similarity is low as such edits were unseen during training. We briefly discuss more limitations in the following section.
6 Conclusions
In this work, we studied the task of motion editing from language instructions. Given the scarcity of training data, we introduced a new dataset MotionFix, collected in a semi-automatic manner. We exploit motion retrieval models to obtain “edit-ready” motion pairs which we annotate with language labels. We design a conditional diffusion model TMED that benefits from this training data, and is able to generate edited motions that follow the source motion and the edit text. We show both quantitatively and qualitatively that our model outperforms all baselines. We hope that our dataset and findings will assist the research community and pave the way for exploring this new task.
Limitations. Our approach comes with limitations. Assuming two TMR-similar motions being an editing distance apart is not always accurate but serves as a good starting point. Another limitation with the data collection is that we constrain the motions to be up to 5 seconds since longer motions tend to produce many dissimilar pairs. In terms of the model performance, TMED exhibits difficulty generalizing to unseen or complex edit texts and maintaining faithfulness to the source motion. Moreover, while our model can be used iteratively, we do not explore this capability in this paper and leave for future work.
Acknowledgments. The authors would like to thank Benjamin Pellkofer for IT support, Tsvetelina Alexiadis for guidance in data collection and perceptual studies, Arina Kuznetcova, Asuka Bertler, Claudia Gallatz, Suraj Bhor, Tithi Rakshit, Taylor McConnell, Tomasz Niewiadomski for data annotation, Lea Müller and Mathis Petrovich for helpful discussions, and Yuliang Xiu for proofreading. GV acknowledges the ANR project CorVis ANR-21-CE23-0003-01.
Disclosure: MJB has received research gift funds from Adobe, Intel, Nvidia, Meta/Facebook, and Amazon. MJB has financial interests in Amazon and Meshcapade GmbH. While MJB is a co-founder and Chief Scientist at Meshcapade, his research in this project was performed solely at, and funded solely by, the Max Planck Society.
@@ Appendix A MotionFix Statistics
We provide additional information about our MotionFix dataset described in Section 3 of the main paper. Since our task is about editing rather than describing motion, it is important that the edits in our dataset contain phraseology of that kind. To confirm this, in Figure A.1 we present qualitative and quantitative visualizations of the word statistics or our dataset’s edit texts. Aside from words that describe body parts, we observe that many of the popular words have an inherently comparative nature. For example, words such as “keep”, “instead”, “slower”, “faster”, “closer”, “earlier” are frequently used as shown in the word cloud visualization. Additionally, in Table A.1 we provide statistics related to the annotated text describing a motion edit. Notice that MotionFix contains a wide spectrum of text descriptions of motion edits ranging from single word commands (e.g., “slower”) to larger paragraphs that describe edits of multiple body parts at different time segments of the motion. Moreover, MotionFix contains 3454, 3171 unique source and target motions, respectively.
@@ Appendix B Additional Quantitative Evaluations
Total #triplets 4771
#Unique texts 4289
#Unique words 1280
Avg #words per text 8.7
@@
Max #words per text 34
Table A.1: Statistics of the MotionFix textual data: There are relatively low number of duplicate texts (given $4771$ triplets and $4289$ unique texts). The vocabulary is diverse ($1280$ unique words) and the average number of words per text ($8.7$) has a good trade-off between conciseness and expressiveness.
Figure A.1: Word frequencies in the MotionFix dataset: On the left, we display a word cloud for the text annotations in in the dataset. Most frequent words appear in larger fonts. Examples of such words are ‘hand’, ‘arm’ referring to body parts, ‘instead’ referring to the source motion, ‘higher’, ‘lower’, ‘opposite’, ‘slower’ referring to spatial, directional or speed edits. On the right, we show the histogram of the $30$ most frequent words in the data.
@@
GT n/a n/a 71.99 91.51 97.45 1.58 25.88 38.61 45.26 26.26
MDM HumanML3D 1.98 4.24 5.66 196.05 1.41 3.54 4.81 211.13
MDMS HumanML3D ✓, init 1.98 3.96 5.52 190.27 2.83 4.53 5.52 196.08
MDM-BPS HumanML3D ✓, init&BP 13.72 24.05 29.28 64.40 47.24 55.87 60.54 40.89
MDM-BP HumanML3D ✓, BP 13.72 22.21 28.85 68.97 46.82 55.02 59.69 44.96
TMED MotionFix ✓, condition 17.11 26.73 34.51 46.42 36.21 51.34 60.25 14.86
Table A.2: Results on the full gallery (comparison to baselines): We evaluate the same models in Tables 2 of the main paper when retrieving among the full test set of 717 samples (as opposed to a random subset of 32). While the retrieval metrics are lower –as expected– due to a larger gallery size, the same conclusions still hold.
@@
GT 97.91 100.0 100.0 1.08 26.36 38.35 46.58 24.96
10% 4.53 7.92 11.03 124.6 8.35 12.87 17.82 81.85
50% 11.60 18.10 24.61 70.29 17.54 27.16 32.53 50.39
100% 17.11 26.73 34.51 46.42 36.21 51.34 60.25 14.86
Table A.3: Results on the full gallery (training data size): Our conclusions from Table 3 of the main paper are confirmed, as we observe significant performance improvement when more portion of MotionFix training data are used.
Additional metrics. To provide a more comprehensive evaluation, besides our retrieval-based metrics, we compute additional measures: FID of motion features and L2 distance between joint positions. We compute both between the test set of MotionFix and the generated motions. For FID, we use the motion branch from the TMR model for the computation of the retrieval metrics. TMED has an FID of $0.136$ outperforming the two best baselines $0.151$ for MDM-BPS and $0.143$ for MDM-BP. While TMED matches best the distribution of target motions, the baselines perform closely indicating that all models maintain certain degree of realism. Our model also has a lower L2 distance (TMED: 1.12cm, MDM-BPS: 1.25cm, MDM-BP: 1.23cm). Note that coordinate-based metrics are not always reliable given multiple plausible generations due to language ambiguity.
@@ Appendix D GPT-4 based annotation in MDM-BP
For the MDM-BP baseline described in Section 5.2 of the main paper, in order to automatically extract which body parts to edit, we prompt GPT-4 through its publicly available API: https://openai.com/api/. Our approach is similar to the one followed in [3]. We feed the edit text to the LLM and prompt it to provide the body parts that should be edited in order to achieve the given motion edit. We use the same body parts as in [3] and give them to GPT-4 as the list to choose from. We experimented with various prompts and eventually used the following: You will be given an edit text that is supposed to be used to edit a motion. Your task is given the text to determine what are the parts of the motion that should change based on that edit text. The instructions for this task are to choose your answers from the list below: left arm, right arm, left leg, buttocks, waist, right leg, torso, neck Here are some examples of the question and answer pairs for this task: ‘‘Question: What are the body parts that should be edited in the motion if the edit text is: faster?’’ ‘‘Answer: right leg, left leg, buttocks, left arm, right arm, torso, neck’’ ‘‘Question: What are the body parts that should be edited in the motion if the edit text is: do it with the opposite leg?’’ ‘‘Answer: right leg, left leg, buttocks’’ ... ‘‘Question: What are the body parts that should be edited in the motion if the edit text is: stop moving in the end?’’ ‘‘Answer: right leg, left leg, buttocks, left arm, right arm, torso, neck’’ ‘‘Question: What are the body parts that should be edited in the motion if the edit text is: [EDIT TEXT]’’
@@
Table A.4: Example GPT responses: We show several input-output pairs for the MDM-BP baseline that uses automatically extracted body parts from edit text. GPT reasonably identifies the body parts that need to be edited e.g., ‘turn around’ involves the whole body, ‘raise your hand faster and a little higher’ involves both arms as this cannot be inferred purely from text.

### v3

@@ 5.3. Ablations
Figure 4. Guidances of conditions: We illustrate the R@1 performance of TMED for generated-to-target (left) and generated-to-source (right) retrieval benchmarks for $s_{L},s_{M_{S}}\in[1,5]$.
[…]
Guidances
Figure 5. TMED generations: We illustrate several generations from our model with overlaid source (red) and generated (blue) motions. We showcase a variety of test cases ranging from elaborate edits (first example in top left) to short commands (e.g., “mirror”). TMED is able to perform both edits that describe temporal (e.g., “slow down”) or spatial (e.g., “raise your arms higher so it is overhead”) modifications.
[…]
Qualitative results
Guidance hyperparameters. In Figure 4, we present how TMED performs across different guidance values for both conditions. $x$-axis controls the text guidance $s_{L}$, and $y$-axis controls the source motion guidance $s_{M_{S}}$ at test time. We report both generated-to-target (left) and source-to-target (right) R@1 retrieval results. We observe that there needs to be a balance between the two guidance values, and that performances decrease towards the extremes (e.g., top left and bottom right corners of the plots, where only one of the two conditions have higher guidance). This highlights the need to rely on both conditions to perform the task.
Figure 6. Failure cases: We show four failure examples from our model. For each sample, we provide the source motion (red) overlaid both with the generation (blue, left) or the ground-truth target motion (green, right). In the top row, we observe that the model may fail to generate the edited motions when the edit text is detailed and the motions differences are subtle. In the bottom row, although the generated motions follow the edit text, they diverge from the source motions.
[…]
Failure Cases
5.4. Qualitative Results
We display several generations from TMED in Figure 5 to enable qualitative assessment. We observe that our model can perform different types of edits such as the addition of actions (“rotate wrists instead of stretching like yawn”), temporal edits in a motion (“get up a bit earlier”), speed edits (“slow down”) and combinations of these. We refer to our supplementary video for dynamic visualizations, which may be easier to interpret.
In Figure 6, we further provide examples of failures cases from TMED. In the top row, we analyze cases with long edit texts. The model struggles with complex details and does not “keep the body straight” in the left example, nor follows “bend arms in the elbows” instructions on the right side, while wider legs are correctly edited. In the bottom row, we illustrate examples where the model faithfully follows the edit text, but does not resemble the source motion. In the left generation, the steps are correctly wider, but the movement does not continue to the similar position as the source motion. Finally, on the right, the body is kneeling down faster as instructed, but towards the opposite direction.
We additionally provide a qualitative comparison in Figure 7, between TMED and various baselines. We provide two comparisons for each baseline (top block for MDMS, middle block for MDM-BPS, and bottom block for MDM-B).
@@ 5.4. Qualitative Results
Finally, we illustrate generations from our strongest baseline MDM-BP in the last two rows. Both generations involve all parts of the body (e.g., “slow down”) making it hard to follow the source motion. In comparison, our model faithfully performs most of the edits. The disadvantage of TMED, on the other hand, might be the generalization to motion pairs where the TMR similarity is low as such edits were unseen during training. We briefly discuss more limitations in the following.
Figure 7. Qualitative comparisons with baselines: We provide example results by comparing TMED against the baselines on the MF test set: MDMS (top), MDM-BPS (middle) and MDM-BP (bottom). First two columns show the source (red) and ground-truth target (green) motions. Third column is reserved for baselines, the last column for TMED. Generations are denoted in blue.
[…]
Comparison with baselines
6. Conclusions
In this work, we studied the task of motion editing from language instructions. Given the scarcity of training data, we introduced a new dataset MotionFix, collected in a semi-automatic manner. We exploit motion retrieval models to obtain “edit-ready” motion pairs which we annotate with language labels. We design a conditional diffusion model TMED that is trained on MotionFix, and generates edited motions that follow the source motion and the edit text. We show both quantitatively and qualitatively that our model outperforms all baselines. We hope that our dataset and findings will assist the research community and pave the way for exploring this new task.
Limitations. Our approach comes with limitations. Assuming two TMR-similar motions being an editing distance apart is not always accurate but serves as a good starting point. Furthermore, in our data collection, we constrain the motions to be up to 5 seconds since longer motions to produce many dissimilar pairs. Regarding model performance, TMED exhibits difficulty generalizing to unseen or complex edit texts and maintaining faithfulness to the source motion. Moreover, while our model can be used iteratively, we do not explore this capability in this paper and leave for future work.
Acknowledgements.
The authors would like to thank Benjamin Pellkofer for building the data exploration website, Tsvetelina Alexiadis for guidance in data collection and perceptual studies, Arina Kuznetcova, Asuka Bertler, Claudia Gallatz, Suraj Bhor, Tithi Rakshit, Taylor McConnell, Tomasz Niewiadomski for data annotation, Lea Müller and Mathis Petrovich for helpful discussions, Yuliang Xiu for proofreading and Peter Kulits for the support and seatmating. GV acknowledges the ANR project CorVis ANR-21-CE23-0003-01. Disclosure: https://files.is.tue.mpg.de/black/CoI_CVPR_2024.txt
@@ Acknowledgements.
Supplementary video. Along with this document, we provide a video on our webpage, which includes visualizations of motions from the MotionFix dataset and comparisons with baselines. In specific: (i) We first briefly describe our goal. (ii) We then show examples from the dataset we collected for different types of edits. (iii) We provide additional qualitative results showing the capability of our method to generate different types of edits. (iv) We provide qualitative comparisons against baselines, highlighting their main problems and how our method overcomes them.
@@ Appendix A MotionFix Statistics
We provide additional information about our MotionFix dataset described in Section 3 of the main paper. Since our task is about editing rather than describing motion, it is important that the edits in our dataset contain phraseology of that kind. To confirm this, in Figure A.1 we present qualitative and quantitative visualizations of the word statistics or our dataset’s edit texts. Aside from words that describe body parts, we observe that many of the popular words have an inherently comparative nature. For example, words such as “keep”, “instead”, “slower”, “faster”, “closer”, “earlier”, “right”, “left”, are frequently used as shown in the word cloud visualization. Additionally, in Table A.1 we provide statistics related to the annotated text describing a motion edit. Notice that MotionFix contains a wide spectrum of text descriptions of motion edits ranging from single word commands (e.g., “slower”) to larger paragraphs that describe edits of multiple body parts at different time segments of the motion. Moreover, MotionFix contains 4649, 4165 unique source and target motions, respectively.
@@ Appendix B Additional Quantitative Evaluations
Total #triplets 6730
#Unique texts 5992
#Unique words 1479
Avg #words per text 8.5
@@
Max #words per text 43
Table A.1. Statistics of the MotionFix textual data: There are relatively low number of duplicate texts (given $6730$ triplets and $5992$ unique texts). The vocabulary is diverse ($1479$ unique words) and the average number of words per text ($8.46$) has a good trade-off between conciseness and expressiveness.
Figure A.1. Word frequencies in the MotionFix dataset: On the left, we display a word cloud for the text annotations in in the dataset. Most frequent words appear in larger fonts. Examples of such words are ‘hand’, ‘arm’ referring to body parts, ‘instead’ referring to the source motion, ‘higher’, ‘lower’, ‘opposite’, ‘slower’ referring to spatial, directional or speed edits. On the right, we show the histogram of the $30$ most frequent words in the data.
@@
GT n/a n/a 64.36 88.75 95.56 1.74 20.83 33.66 40.47 33.13
MDM HumanML3D 0.00 0.30 0.49 476.65 0.10 0.10 0.10 486.26
MDMS HumanML3D ✓, init 0.00 0.10 0.10 477.68 0.00 0.10 0.30 485.02
MDM-BPS HumanML3D ✓, init&BP 8.39 14.81 18.36 181.38 30.11 36.72 40.77 107.11
MDM-BP HumanML3D ✓, BP 8.69 14.71 18.36 180.99 30.21 36.82 40.47 106.05
TMED MotionFix ✓, condition 14.51 21.72 28.73 56.63 22.41 34.45 40.57 31.42
Table A.2. Results on the MotionFix benchmark using the whole test set as a gallery: We evaluate the models in Table 2 of the main paper on the full test set of $1013$ samples (as opposed to a random subset of 32). While the retrieval metrics are unsurprisingly lower due to a larger gallery size, the conclusions hold.
@@
GT 64.36 88.75 95.56 1.74 20.83 33.66 40.47 33.13
10% 2.57 3.75 5.03 259.73 3.75 5.33 7.40 213.17
50% 8.00 14.71 18.36 104.85 13.03 20.73 25.07 77.08
100% 14.51 21.72 28.73 56.63 22.41 34.45 40.57 31.42
Table A.3. Text-based motion editing benchmark on MotionFix test set with different training data sizes. We observe that the performance increases significantly when more data are used during training.
Additional metrics. To provide a more comprehensive evaluation, besides our retrieval-based metrics, we compute additional measures: FID of motion features and L2 distance between joint positions. We compute both between the test set of MotionFix and the generated motions. For FID, we use the motion branch from the TMR model for the computation of the retrieval metrics. TMED has an FID of $0.129$ outperforming the two best baselines $0.152$ for MDM-BPS and $0.145$ for MDM-BP. While TMED matches best the distribution of target motions, the baselines perform closely indicating that all models maintain certain degree of realism. Our model also has a lower L2 distance (TMED: 1.10cm, MDM-BPS: 1.26cm, MDM-BP: 1.22cm). Note that coordinate-based metrics are not always reliable given multiple plausible generations due to language ambiguity.
@@ Appendix D GPT-4 based annotation in MDM-BP
For the MDM-BP baseline described in Section 5.2 of the main paper, in order to automatically extract which body parts to edit, we prompt GPT-4 through its publicly available API: https://openai.com/api/. Our approach is similar to the one followed in (Athanasiou et al., 2023). We feed the edit text to the LLM and prompt it to provide the body parts that should be edited in order to achieve the given motion edit. We use the same body parts as in (Athanasiou et al., 2023) and give them to GPT-4 as the list to choose from. We experimented with various prompts and eventually used the following: You will be given an edit text that is supposed to be used to edit a motion. Your task is given the text to determine what are the parts of the motion that should change based on that edit text. The instructions for this task are to choose your answers from the list below: left arm, right arm, left leg, buttocks, waist, right leg, torso, neck Here are some examples of the question and answer pairs for this task: ‘‘Question: What are the body parts that should be edited in the motion if the edit text is: faster?’’ ‘‘Answer: right leg, left leg, buttocks, left arm, right arm, torso, neck’’ ‘‘Question: What are the body parts that should be edited in the motion if the edit text is: do it with the opposite leg?’’ ‘‘Answer: right leg, left leg, buttocks’’ ... ‘‘Question: What are the body parts that should be edited in the motion if the edit text is: stop moving in the end?’’ ‘‘Answer: right leg, left leg, buttocks, left arm, right arm, torso, neck’’ ‘‘Question: What are the body parts that should be edited in the motion if the edit text is: [EDIT TEXT]’’
@@
Table A.4. Example GPT responses: We show several input-output pairs for the MDM-BP baseline that uses automatically extracted body parts from edit text. GPT reasonably identifies the body parts that need to be edited e.g., ‘turn around’ involves the whole body, ‘raise your hand faster and a little higher’ involves both arms as this cannot be inferred purely from text.